When two numbers are added together, the result is 45. Twice the greater number is six more than five times the lesser number. What are the numbers?
Answer by mananth(16949)

let x be the smaller number'
y the bigger number
x+y=45......................1
2y=5x+6
2y-5x=6......................2
multiply equation 1 by 5 and add to 2
5x+5y=225
add to 2
7y=231
/7
y =33
...
plug value of y in 1
x+21=45
x=24
24 & 33
...
m.ananth@hotmail.ca